In racing over a given distance d at a uniform speed A can beat B by 30 M  ,B can beat C  by 20  M and A can beat C by 48 m . Find D in metres?  With Explanation

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
Let speed of A,B,C and D are a, b, c, d respectively.
 
D/a = D – 30/b           (i)
 
D/b = D – 20/c                (ii)
 
D/a = D – 48/c         (iii)
 
On solving
 
D = 300 metre
 
Hope it helps
 
Regards
Arun